OpenStudy (anonymous):

1/cot²40° -1/cos²40° =

OpenStudy (anonymous):

Expand, Lcm and simplify.

OpenStudy (anonymous):

\[\cot^2(x)=\frac{\cos^2(x)}{\sin^2(x)}\] so \[\frac{1}{\cot^2(x)}=\frac{\sin^2(x)}{\cos^2(x)}\]

OpenStudy (anonymous):

\[\frac{\sin^2(x)}{\cos^2(x)}-\frac{1}{\cos^2(x)}=\frac{\sin^2(x)-1}{\cos^2(x)}\] \[=\frac{\cos^2(x)}{\cos^2(x)}=1\]